### v3

@@ Abstract
3D Gaussian Splatting has emerged as an alternative 3D representation for novel view synthesis, benefiting from its high-quality rendering results and real-time rendering speed. However, the 3D Gaussians learned by 3D-GS have ambiguous structures without any geometry constraints. This inherent issue in 3D-GS leads to a rough boundary when segmenting individual objects. To remedy these problems, we propose SAGD, a conceptually simple yet effective boundary-enhanced segmentation pipeline for 3D-GS to improve segmentation accuracy while preserving segmentation speed. Specifically, we introduce a Gaussian Decomposition scheme, which ingeniously utilizes the special structure of 3D Gaussian, finds out, and then decomposes the boundary Gaussians. Moreover, to achieve fast interactive 3D segmentation, we introduce a novel training-free pipeline by lifting a 2D foundation model to 3D-GS. Extensive experiments demonstrate that our approach achieves high-quality 3D segmentation without rough boundary issues, which can be easily applied to other scene editing tasks. Our code is public available at https://github.com/XuHu0529/SAGS.
Keywords:
3D Gaussian Splatting 3D Segmentation Boundary Issues
1 Introduction
@@
2 Related Work
2.1 3D Gaussian Splitting
@@ 2 Related Work
3D Gaussian Splatting [12] is a technique for scene reconstruction, which has been shown to be an alternative 3D representation as NeRF, benefiting from both its high-quality rendering results and real-time rendering speed. Recent research on this technique contains applying 3D Gaussian in dynamic scene and combining it with diffusion model to achieve 3D generation. Dynamic 3D Gaussians [19] extends Gaussian Splatting to dynamic scenes via tracking 3D objects modeled as a set of 3D Gaussians. DreamGaussian [31], GaussianDreamer [35] and GSGEN [6] combine 3D Gaussians with diffusion models to generate high-quality 3D assets.
2.2 Segmentation in Radiance Fields
@@
Neural Radiance Fields (NeRFs) are a popular way of representing 3D scenes implicitly with neural networks. Many researchers have explored how to segment objects in 3D using NeRFs, for various applications such as novel view synthesis, semantic segmentation, 3D inpainting, and language grounding. Some methods, such as Semantic-NeRF [36], NVOS [28], and SA3D [4], use different types of inputs to guide the segmentation, such as semantic labels, user scribbles, or 2D masks. Other methods, such as N3F [32], DFF [32], LERF [13], and ISRF [9], learn additional feature fields that are aligned with NeRFs, and use 2D visual features from pre-trained models or language embeddings to query the 3D features. These methods usually require modifying or retraining the original NeRF models or training other specific parameters to obtain the 3D segmentation. However, limited by the representation and rendering speed of NeRF, it is challenging to apply it to more practical applications, such as scene editing, collision analysis, etc. Inspired by the real-time 3D-GS, recent works [3, 34, 37, 16] have been proposed to achieve segmentation in 3D-GS in a learnable way. SAGA [3] and Featrue 3DGS [37] distill the knowledge embedded in the SAM decoder into the feature field of 3D-GS, and Gaussian-Grouping [34] supervises the Identity Encodings during the differentiable rendering by leveraging the 2D mask predictions by SAM. However, no research has been proposed to solve the rough boundary issues in segmentation resulting from the inherent properties of 3D Gaussians. In this work, we are the first to propose a novel Gaussian Decomposition scheme to solve this issue, incorporated into our effective training-free segmentation pipeline.
@@
Figure 2: Pipeline of our proposed method. (a) Given a set of clicked points on the $1^{st}$ rendered view, we utilize SAM to generate masks for corresponding objects under every view automatically; (b) For every view, Gaussian Decomposition is first performed to address the issue of boundary roughness and then label propagation is implemented to assign binary labels to each 3D Gaussian; (c) Finally, with assigned 3D labels from all views, we adopt a simple yet effective voting strategy to determine the segmented Gaussians.
3 Method
@@
In this section, we first present the preliminary in 3D Gaussian Splatting in Sec. 3.1 for clear understanding and then define the problem and task in Sec. 3.2. Our basic training-free pipeline consists of Sec. 3.3 and Sec. 3.5. The details of Gaussian Decomposition are described in Sec. 3.4. The pipeline of our method can be seen in Fig. 2.
3.1 Preliminary: 3D Gaussian Splatting
@@
3D Gaussian Splatting (3D GS) [12] is an emerging method for real-time radiance field rendering. It has been proven effective in Novel View Synthesis with high rendering quality as NeRF and real-time rendering speed. 3D GS represents scenes with a set of 3D Gaussians. Specifically, each 3D Gaussian is parameterized by a position $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma$, an opacity value $\alpha$, and spherical harmonics (SH).
@@
3.2 Problem Definition
@@
3.3 Segment 3D Gaussians with 2D Mask
3D Prompts for Multiview Masks Generation. By the definition in Sec. 3.2, users are given the first rendered view to specify the target object. However, only the first view is far from sufficient to segment the target object in 3D space. So here we first generate multiview masks to aid the 3D segmentation. With multiple masks from different views, a 3D object can be segmented by the intersection of the corresponding frustum of these masks.
@@ 3.3 Segment 3D Gaussians with 2D Mask
$\operatorname*{arg\,min}_{\mu}\{d(\mu),d(\mu)>0\mid\mu\in\mathbb{G},\|\mathbf{P}_{0}\mu-\mathbf{p}_{i}^{0}\|_{1}<\epsilon\},$ (2)
@@
For all the 2D prompt points in the first view, we could get a set of 3D prompt by Eq. 2. Then for another view $\mathbf{v}_{i}$, we project these 3D prompts into the 2D plane, resulting in 2D prompts in the view $\mathbf{v}_{i}$. In this way, we obtain 2D prompts in all views, and all masks are obtained by prompting the SAM.
@@
Figure 3: Illustration of the Gaussian Decomposition process.
3.4 Gaussian Decomposition
@@
[…]
@@
[…]
@@
3.5 Multiview Label Voting
So far, every 3D Gaussian including the decomposed one has a list of binary labels $\mathbf{L}_{i}$, using the label assignment in Sec. 3.3. Leveraging the assigned labels, here we adopt a simple yet effective heuristic rule to determine if a 3D Gaussian $\mathbf{g}_{i}$ belongs to the target 3D object. In particular, we first define the confidence score $s_{i}$ of $g_{i}$ as
@@ 3.5 Multiview Label Voting
where $N$ is the number of views. Then we adopt a threshold $\tau$ and regard a Gaussian with a score higher than $\tau$ as positive.
4 Experimets
4.1 Datasets
We choose different datasets to testify our method, including LLFF [22], Mip-NeRF 360 [2], LERF [13], and some test scenes from the 3D Gaussian Splatting [12]. These datasets contain both small indoor objects and large outdoor scenes, which are very complex and challenging. For quantitative experiments, because there is no existing benchmark that can be used in 3D Gaussian space, we use the SPIn-NeRF [23] dataset with 2D ground-truth for evaluation.
4.2 Quantitative results
Table 1: Quantitative results on SPIn-NeRF dataset. “Singe view” denotes projecting the 2D segmentation result to 3D simply, thus we don’t consider its time cost. There is extra time cost (about 10mins) for SAGA to distill 2D knowledge to 3D-GS, while other methods don’t need this process.
@@ 4.2 Quantitative results
Lego 76.0 99.1 74.9 99.2 92.2 99.8 - - 88.4 99.7
Mean 74.1 95.2 90.5 98.8 91.9 98.8 88.0 98.5 89.9 98.7
Time cost - 3-6mins 2-10mins 0.08s-0.9s within 15s
(/object) (+10mins)
We first conduct experiments on the SPIn-NeRF [23] dataset for quantitative analysis. Given a set of images of the scene, we follow the process described in Section 3 to obtain the segmentation of the target object in 3D Gaussian Space. The segmented 3D Gaussians are used to render 2D masks in other views. Finally, we calculate the IoU and Accuracy between these rendered and the ground-truth masks. Results can be seen in Table 1.
In the comparison, it is noteworthy that both the MVSeg [23] and the SA3D [4] require additional parameters and a computation-costly training process. While SAGA [3] shows nearly one-thousandth of the time compared with them, it needs extra training time (10 minutes per scene) to distill the knowledge embedded in the SAM decoder into the feature field. By contrast, the “Single view” [4] refers to mapping the 2D masks to the 3D space based on the corresponding depth information, which does not need additional training process. In this sense, our method is the same as the “Single view” that does not incorporate any additional training or model parameters. Results show that our approach can achieve comparable segmentation quality to the MVSeg and SA3D methods and better results than SAGA. The time cost of ours is much less than MVSeg and SA3D. Considering the training time of SAGA, the average segmentation cost of a single object is basically the same as our method. In some 360 outdoor scenes, such as “Truck”, our approach even outperforms the SA3D and the MVSeg. When compared with the training-free method “Single view”, our approach achieves a significant promotion of +15.8% IoU and +3.5% Acc. These results demonstrate our approach is very efficient in obtaining high-quality segmentation masks.
Table 2: Quantitative results with text prompts on four scenes of SPIn-NeRF dataset.

### v4

@@ Abstract
3D Gaussian Splatting has emerged as an alternative 3D representation for novel view synthesis, benefiting from its high-quality rendering results and real-time rendering speed. However, the 3D Gaussians learned by 3D-GS have ambiguous structures without any geometry constraints. This inherent issue in 3D-GS leads to a rough boundary when segmenting individual objects. To remedy these problems, we propose SAGD, a conceptually simple yet effective boundary-enhanced segmentation pipeline for 3D-GS to improve segmentation accuracy while preserving segmentation speed. Specifically, we introduce a Gaussian Decomposition scheme, which ingeniously utilizes the special structure of 3D Gaussian, finds out, and then decomposes the boundary Gaussians. Moreover, to achieve fast interactive 3D segmentation, we introduce a novel training-free pipeline by lifting a 2D foundation model to 3D-GS. Extensive experiments demonstrate that our approach achieves high-quality 3D segmentation without rough boundary issues, which can be easily applied to other scene editing tasks. Our code is publicly available at https://github.com/XuHu0529/SAGS.
Index Terms:
3D gaussian splatting, 3D segmentation, boundary issues.
I Introduction
@@
II Related Work
II-A Radiance Fields and 3D Gaussian Splatting
Novel view synthesis (NVS) involves rendering unseen viewpoints of a scene from a given set of images. One popular approach is Neural Radiance Fields (NeRF), which uses a Multilayer Perceptron (MLP) to represent 3D scenes and map from a 3D coordinate to properties of the scene at the corresponding location. It leverages the differentiable volume rendering technique to translate a 3D scene’s continuous representation into 2D images. Several works have been proposed to enhance NeRF’s performance by addressing aspects such as speed [78, 79, 10, 59, 24, 25, 72, 94, 60] and adapting it to other tasks [77, 74, 85, 105]. Though some breakthroughs have been achieved, the reliance on low-efficient volume rendering still hinders real-time rendering.
@@ II-A Radiance Fields and 3D Gaussian Splatting
Recently, 3D Gaussian Splatting [35] has been proposed as a new technique for novel view synthesis. It leverages an ensemble of anisotropic 3D Gaussian splats to represent the scene and employs differentiable splatting for rendering. 3D Gaussian Splatting has been shown to be an alternative 3D representation of NeRF, benefiting from both its high-quality rendering results and real-time rendering speed. Recent research on this technique involves applying 3D Gaussian Splatting to large-scale scene reconstruction [49, 50], the dynamic scenes [52, 87, 46, 112, 98, 1, 51, 75, 31, 45, 95, 41, 80], 3D object and scene generation [81, 93, 16, 89, 90, 111, 43, 103, 109], surface reconstruction [30, 97, 20, 27] and other tasks [9, 86, 44, 75].
II-B 2D and 3D Perception
Detection and segmentation are fundamental computer vision tasks. Numerous studies [68, 8, 5, 66, 67, 106, 64, 101, 65, 102] have deeply explored various sub-fields. Especially, Grounding-DINO [47] enhances the grounding results by introducing the language instructions. Traditionally, the segmentation includes three major tasks: semantic [12, 14, 13, 15, 84], instance [28, 48, 4], and panoptic [37] segmentation. Due to the operation consistency of the above tasks at the pixel level, many studies have tried to use a unified framework, such as K-net [104], MaskFormer [18], and Mask2Former [17]. A significant breakthrough in 2D segmentation is the Segment Anything Model (SAM) [38]. SAM seeks to unify the 2D segmentation task using a prompt-based segmentation approach. Its introduction has sparked a new wave of research, with many studies focused on enhancing its functionality. These improvements include efficient fine-tuning techniques [70, 71, 34] and distillation-based acceleration methods [107, 99]. Additionally, SAM has been adapted for various fields, such as medical image analysis [53, 54, 88, 21], concealed object detection [32, 82], image editing [96], remote sensing [11], and 3D bird’s eye view (BEV) sensing [100].
II-C Segmentation in Radiance Fields
@@
Neural Radiance Fields (NeRFs) are a popular way of implicitly representing 3D scenes with neural networks. Many researchers have explored how to segment objects in 3D using NeRFs for various applications such as novel view synthesis, semantic segmentation, 3D inpainting, and language grounding. Some methods, such as Semantic-NeRF [108], NVOS [73], and SA3D [7], use different types of inputs to guide the segmentation, such as semantic labels, user scribbles, or 2D masks. Other methods, such as N3F [83], DFF [83], LERF [36], and ISRF [26], learn additional feature fields that are aligned with NeRFs, and use 2D visual features from pre-trained models or language embeddings to query the 3D features. These methods usually require modifying or retraining the original NeRF models or training other specific parameters to obtain the 3D segmentation. However, limited by the representation and rendering speed of NeRF, it is challenging to apply it to more practical applications, such as scene editing, collision analysis, etc. Inspired by the real-time 3D-GS, recent works [6, 92, 110, 40] have been proposed to achieve segmentation in 3D-GS in a learnable way. SAGA [6] and Featrue 3DGS [110] distill the knowledge embedded in the SAM decoder into the feature field of 3D-GS, and Gaussian-Grouping [92] supervises the Identity Encodings during the differentiable rendering by leveraging the 2D mask predictions by SAM. However, no research has been proposed to solve the rough boundary issues in segmentation resulting from the inherent properties of 3D Gaussians. In this work, we are the first to propose a novel Gaussian Decomposition scheme to solve this issue, incorporated into our effective training-free segmentation pipeline.
@@
Figure 2: Pipeline of our proposed method. (a) Given a set of clicked points on the $1^{st}$ reference view, we utilize SAM to generate masks for corresponding objects under every view automatically; (b) For every view, Gaussian Decomposition is first performed to address the issue of boundary roughness and then label propagation is implemented to assign binary labels to each 3D Gaussian; (c) Finally, with assigned 3D labels from all views, we adopt a simple yet effective voting strategy to determine the segmented Gaussians.
III Method
@@
In this section, we first present the preliminary in 3D Gaussian Splatting and Segment Anything Model (SAM) in Sec. III-A for clear understanding and then define the problem and task in Sec. III-B. Our basic training-free pipeline consists of Sec. III-C and Sec. III-E. The details of Gaussian Decomposition are described in Sec. III-D. The pipeline of our method can be seen in Fig. 2.
III-A Preliminary
3D Gaussian Splatting
@@ III-A Preliminary
3D Gaussian Splatting (3D GS) [35] is an emerging method for real-time radiance field rendering. It has been proven effective in Novel View Synthesis with high rendering quality as NeRF and real-time rendering speed. 3D GS represents scenes with a set of 3D Gaussians. Specifically, each 3D Gaussian is parameterized by a position $\mu\in\mathbb{R}^{3}$, a covariance matrix $\Sigma$ consisting of a scaling factor $s\in\mathbb{R}^{3}$, and a rotation quaternion $q\in\mathbb{R}^{4}$, an opacity value $\alpha$, and spherical harmonics (SH). Each 3D Gaussian is characterized by:
$G(x)=\frac{1}{\left(2\pi\right)^{3/2}\left|\Sigma\right|^{1/2}}e^{-\frac{1}{2}{\left(x-\mu\right)}^{T}\Sigma^{-1}\left(x-\mu\right)}$ (1)
@@
Segment Anything Model (SAM)
SAM[38] takes an image $I$ and a set of prompts $P$ as input to output the corresponding segmentation mask $M_{SAM}$:
$M_{SAM}=SAM(I,P)$ (3)
where the prompts $P$ can be points, bounding boxes, or texts.
III-B Problem Definition
@@
III-C Segment 3D Gaussians with 2D Mask
3D Prompts for Multiview Masks Generation. By the definition in Sec. III-B, users are given the first rendered view to specify the target object. However, only the first view is far from sufficient to segment the target object in 3D space. So here we first generate multiview masks to aid the 3D segmentation. With multiple masks from different views, a 3D object can be segmented by the intersection of the corresponding frustum of these masks.
@@ III-C Segment 3D Gaussians with 2D Mask
$\arg\min_{\mu}\{d(\mu),d(\mu)>0\mid\mu\in\mathbb{G},\|\mathbf{P}_{0}\mu-\mathbf{p}_{i}^{0}\|_{1}<\epsilon\},$ (4)
@@
For all the 2D prompt points in the first view, we could get a set of 3D prompts by Eq. 4. Then for another view $\mathbf{v}_{i}$, we project these 3D prompts into the 2D plane, resulting in 2D prompts in the view $\mathbf{v}_{i}$. In this way, we obtain 2D prompts in all views, and all masks are obtained by prompting the SAM.
@@
Figure 3: Illustration of the Gaussian Decomposition process. It involves two basic steps: first, to find out the boundary Gaussians and then decompose these Gaussians.
III-D Gaussian Decomposition
@@
To address this issue, we propose the Gaussian Decomposition to mitigate the boundary roughness while maintaining the 3D structures as complete as possible. Fig. 3 illustrates the basic idea of Gaussian decomposition. It has two basic steps to achieve the Gaussian decomposition:
(1) For each 3D Gaussian, we obtain the corresponding 2D Gaussian by projection and mark it as a boundary Gaussian if one of its long-axis endpoints is outside of the 2D mask. Then, as shown in Fig. 3, we denote the two ends of the long axis of a 2D Gaussian as $A^{*}$ and $B^{*}$, and the intersection of $A^{*}B^{*}$ and mask boundary is $O^{*}$. $A$, $B$, and $O$ are the corresponding 3D points in the long axis of the 3D Gaussian. Assuming $A^{*}$ is in the mask while $B^{*}$ is out of the mask, we define
@@
Let $\mathbf{g}$ denote a 3D Gaussian across the boundary. Its scale in the long axis and the 3D center are defined as $s$ and $\mathbf{\mu}$, respectively. We have
@@
III-E Multiview Label Voting
So far, every 3D Gaussian including the decomposed one has a list of binary labels $\mathbf{L}_{i}$, using the label assignment in Sec. III-C. Leveraging the assigned labels, here we adopt a simple yet effective heuristic rule to determine if a 3D Gaussian $\mathbf{g}_{i}$ belongs to the target 3D object. In particular, we first define the confidence score $s_{i}$ of $g_{i}$ as
@@ III-E Multiview Label Voting
where $N$ is the number of views. First, $s_{i}>0.5$ is performed to vote out the object 3D Gaussian. Then, to reduce the background bias due to occlusion, we adopt a threshold $\tau$ and an Object-ID $\mathbf{O}_{i}$ for each 3D Gaussian $\mathbf{g}_{i}$ is determined by:
$\mathbf{O}_{i}=\begin{cases}1&\text{if }s_{i}>\tau,\\ 0&\text{if }s_{i}<\tau,\end{cases}$ (12)
III-F Multi-object Segmentation
There are numerous objects in the 3D scenes. Our method can also support segmenting all objects simultaneously by merely extending the binary masks to multi-label masks. Specifically, the element $\mathbf{L}_{ij}$ in the matrix $\mathbf{L}$ is extended from binary values to multiple values, where $\mathbf{L}_{ij}\in\{0,1,...,C-1\}$ and $C$ is the number of objects. The similar multi-view label voting strategy will determine an Object-ID $\mathbf{O}_{i}$ for each 3D Gaussian $\mathbf{g}_{i}$:
$\mathbf{O}_{i}=Mode(\mathbf{L}_{ij}),j\in\{0,1,...,N-1\},$ (13)
where $Mode(.)$ is the function of finding the element that appears the most, and $N$ is the number of views.
IV Experimets
IV-A Datasets
We choose different datasets to testify our method, including LLFF [57], Mip-NeRF 360 [3], LERF [36], and some test scenes from the 3D Gaussian Splatting [35]. These datasets contain both small indoor objects and large outdoor scenes, which are very complex and challenging. For quantitative experiments, because there is no existing benchmark that can be used in 3D Gaussian space, we use the SPIn-NeRF [58] dataset with 2D ground truth for evaluation.
IV-B Implementation Details
As the implementation of 3D Gaussian Splatting [35] for each scene, we follow the official code with default parameters, and each scene is trained with 30000 iterations. For generating single object segmentation masks, we don’t limit the number of clicked points on a single reference view for SAM. Binary masks of other views can be automatically generated via the proposed method. This is reasonable since users can refine their input prompts to help SAM generate a 2D mask as accurately as possible from the reference view. For scene segmentation to generate multi-label masks, we first employ SAM to produce instance masks for individual views under automatic segmentation mode. To ensure 2D mask consistency across views, we use a pre-trained zero-shot tracker [19, 92] to propagate and associate masks. For the selection of hyper-parameters, we only have one to control, which is the confidence score threshold $\tau$ in the Multiview Label Voting method III-E. By default, we set the value to 0.7 in all experiments. Also, in practical use, users can manually set this value according to the complexity of different scenes. As for the number of views used in experiments, we follow SA3D [7] to select all view images of each scene to finish our segmentation process.
IV-C Quantitative results
TABLE I: Quantitative results on SPIn-NeRF dataset. Singe view denotes projecting the 2D segmentation result to 3D simply.
@@ IV-C Quantitative results
Lego 76.0 99.1 74.9 99.2 92.2 99.8 - - 90.2 99.7
Mean 74.1 95.2 90.5 98.8 91.9 98.8 88.0 98.5 90.0 98.7
Point-Guided Segmentation: We first conduct experiments on the SPIn-NeRF [58] dataset for quantitative analysis. Given a set of images of the scene, we follow the process described in Section III to obtain the segmentation of the target object in 3D Gaussian Space. The segmented 3D Gaussians are used to render 2D masks in other views. Finally, we calculate the IoU and Accuracy between these rendered and the ground-truth masks. Results can be seen in Table I.
In the comparison, it is noteworthy that both the MVSeg [58] and the SA3D [7] require additional parameters and a computation-costly training process. While SAGA [6] shows nearly one-thousandth of the time compared with them, it needs extra training time (10 minutes per scene) to distill the knowledge embedded in the SAM decoder into the feature field. By contrast, the “Single view” [7] refers to mapping the 2D masks to the 3D space based on the corresponding depth information, which does not need additional training process. In this sense, our method is the same as the “Single view” that does not incorporate any additional training or model parameters. Results show that our approach can achieve comparable segmentation quality to the MVSeg and SA3D methods and better results than SAGA. The time cost of ours is much less than MVSeg and SA3D. Considering the training time of SAGA, the average segmentation cost of a single object is basically the same as our method. In some 360 outdoor scenes, such as “Truck”, our approach even outperforms the SA3D and the MVSeg. When compared with the training-free method “Single view”, our approach achieves a significant promotion of +15.9% IoU and +3.5% Acc. These results demonstrate our approach is very efficient in obtaining high-quality segmentation masks.
TABLE II: Quantitative results with text prompts on four scenes of SPIn-NeRF dataset.
@@
Text-Guided Segmentation: We replace the point prompts with text prompts corresponding to the objects. Given a text input, we prompt the Grounding DINO [47] to generate target bounding boxes, which serve as input prompts for the SAM to obtain 2D segmentation masks. Then, following the segmentation process in Section III, we can transform the 2D masks into 3D Gaussian masks corresponding to the text input. Similar to the above evaluation process, we also calculate the IoU and Acc between 2D rendered masks and given ground-truth masks. Quantitative results are shown in Table II. For both room and truck scenes, the IoU and Acc values can achieve a similar level as using clicked points as input prompts. These results demonstrate our method can combine multi-modal prompts as input.
IV-D Boundary Segmentation Analysis
The main contribution of our work is to address the boundary issues of segmentation in 3DGS, as mentioned in Sec. I and Fig. 1. To validate the effectiveness of our work, we evaluate metrics, including IoU, AP, and F1-score, on region nearby boundaries. Specifically, we extract parts of 3-pixel width alongside boundaries of ground-truth masks and compare results with previous work SAGA [6]. Experiments are conducted on Scene pinecone and fortress from the LLFF [57] dataset. Better results of ours are indicated in Tab. III. Benefiting from our proposed Gaussian Decomposition for ambiguous geometry around boundaries, we can achieve more complete and accurate boundary segmentation results.